OpenStudy (anonymous):

solve the equation:3/(x+1) -1/x=1/2

OpenStudy (mertsj):

Multiply each term by the common denominator which is 2x(x+1)

OpenStudy (blurbendy):

Common denominator is x(x + 1) So After simplifying you should get (2x -1) / x(x+1) = 1/2 You can cross multiply here and get 2(2x -1) = x(x+1) Now move everything to left side 2(2x -1) -x(x+1) = 0 Expand and get -x^2 +3x -2 = 0 -((x - 2)(x -1) = 0 x = 2 x = 1

OpenStudy (anonymous):

are you sure?

OpenStudy (anonymous):

because after i submit them i cant do it again^.^

OpenStudy (blurbendy):

Yup, see what happens when you put 2 in for x into the original equation. Then see what happens when you put 1 in for x. You'll always get the right answer. I'm 100 percent positive.

OpenStudy (anonymous):

OK:)